OpenStudy (anonymous):

Solve the equation by completing the square. x^2 + 6x − 14 = 0 x = (smaller value) x = (larger value)

OpenStudy (cruffo):

\(x^2 + 6x = 14\) Take half the bx term , square it, and add the result to both sides: \(\frac{1}{2}(6) = 3\) and \(3^2 = 9\). So \(x^2 + 6x +9 = 14 +9\) Now factor the left and simplify the right, \((x+3)^2 = 23\)

OpenStudy (anonymous):

x^2+6x+9 = 23?

OpenStudy (radar):